OpenStudy (anonymous):

If a golf ball is dropped from a tall building, the distance it has fallen in feet is given by 16t2, where t is the time in seconds since the ball was dropped. How far has the ball fallen after 4 seconds?

OpenStudy (anonymous):

Hi, all you need to do is substitute t with 4 in the equation: f(x)=16(t)^2. So: f(4)=16(4)^2. = 16(16) = 256 feet. It is that simple.

OpenStudy (anonymous):

thank you so much I have the answer earlier but im not sure about it.

OpenStudy (anonymous):

No problem. Ask anytime.

OpenStudy (anonymous):

u guys help me a lot :-)

OpenStudy (anonymous):

Haha, we're just freelancers ;)

OpenStudy (anonymous):

gees lucky you